OpenStudy (anonymous):

please help!! 1 medal and fan :) TQ congruent to QR. Find m

Directrix (directrix):

The triangles are congruent by the HL Theorem. So corresponding angles are congruent. Your First Task: Because the two angle with vertex S are congruent, solve this equation for x: 2x + 13 = 21 What is x? @knowel

OpenStudy (anonymous):

4

OpenStudy (anonymous):

thats it right?

Directrix (directrix):

What is 2*4 + 13 = ?

OpenStudy (anonymous):

21

Directrix (directrix):

The sum of the interior angle measures of triangle SQT is 180. So, solve this equation for y: y = 180 - 21 - 90 y =? ---> That is the value of m<SQT. So, what is y?

Directrix (directrix):

180 - 21 - 90 = ? @knowel

OpenStudy (anonymous):

69

Directrix (directrix):

That is what I got.

OpenStudy (anonymous):

thanks